# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v2

@@ Abstract
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. These models outperform six state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. We also empirically show that the tuned models are robust to adaptive jailbreaks and prompt injections.
@@ 1 Introduction
The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries. To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [62, 99, 40, 12, 47, 33, 10, 17]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-97.26% (average: 65.06%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.244, and GPT-4-based SelfDefend even reduces the ASR by 69.69-100% (average: 85.53%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.063. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [34] and find that the pass rate is almost unaffected for GPT-3.5 and slightly decreases by 2.77% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations. For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.05 seconds for GPT-3.5 and 0.40 seconds for GPT-4, respectively.
@@
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063, GPT-4 itself is commonly known to be expensive [3]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [19] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data, which is then used to tune our defense models through LoRA fine-tuning [22] on the publicly available Llama-2-7b model [67]. The details will be introduced in §6.
@@
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with six representative jailbreak defenses: ICD [73], SafeDecoding [82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25], and Llama Guard 2 [66]. The results show that SelfDefend consistently outperforms all other defenses in 20 out of 24 tested attack scenarios (involving eight jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama-2) and reaches the level of GPT-4-based SelfDefend’s defense.
@@
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.27 seconds, with $\Delta d$ in all attack scenarios except for DAN now below 0.1 seconds. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it efficient for potential real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [58], and empirically show that the tuned models are robust to adaptive attacks and prompt injections [44, 42]. Details are available in §7.
@@ 2.2 Jailbreak Attacks
Optimization-based Jailbreak typically updates the adversarial prompt iteratively using gradient-based or search-based methodologies [99, 63, 5, 40, 28]. The pioneering work by GCG [99] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [63] further extended this technique to GCG++ by employing a proxy model to direct the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [5] utilized a simple random search on a suffix to increase the likelihood of hitting the target probability. Unlike optimizing these obviously unreadable suffixes, AutoDAN [40] automatically constructs human-readable jailbreak prompts using a carefully designed hierarchical genetic algorithm.
Generation-based Jailbreak employs language models [54, 16, 12, 47, 53] to generate effective jailbreak prompts that can mislead LLMs into producing restricted content. An intuitive approach is to use an auxiliary LLM to construct candidate prompts through prompt engineering. For example, PAIR [12] fed the response of the target model back to the attacking LLM to adapt the output for deceptive jailbreaks. Mehrotra et al. [47] then refined PAIR’s approach through tree-of-thought reasoning [84]. Moreover, the adversary can train a new LLM specifically to attack the target model. For example, Paulus et al. [53] fine-tuned the Advprompter LLM to generate human-readable suffixes against the target LLM.
@@
Gradient Cuff [23] arXiv:2403.00867 Compare the gradient norm of refusal loss with that in benign queries
Circuit Breaking [98] arXiv:2406.04313 Interrupt the model output harmful content at the internal representations
@@
RPO [96], DPP [81] arXiv:2401.17263 Optimize universal and transferable suffixes that enforce safe outputs
@@
CAT [77] arXiv:2405.15589 Conduct adversarial training on continuous embedding attacks
LED [91] arXiv:2405.18166 Identify safety-critical layers and realign them through model editing
@@ 4 The SelfDefend Framework
After analyzing the pitfalls of existing defenses, we propose a new perspective on defending LLMs against jailbreaks. Our key idea is to deploy a dedicated LLM alongside the target LLM to concurrently detect potential jailbreak queries. This idea is made possible because we found that LLMs do have the capability to protect themselves by identifying harmful portions in user queries.
Overview. Based on this idea, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [8] and the library-based checkpoint from SCLib [74]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries.
Workflow. ① Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. ② $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal prompt or an adversarial prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. By contrast, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful part (via $P_{direct}$) or intention (via $P_{intent}$). ③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{defense}$ is triggered to release its token-by-token response. Otherwise, when the shadow stack detects a harmful portion (prompt/intention), SelfDefend would respond with a template that refuses to answer, i.e., “I can’t fulfill your query because your [harmful portion] violated our safety policy.” It is noted that “[harmful portion]” is replaced with the recognized portion from $LLM_{defense}$.
@@ 5 An Empirical Measurement
Besides initializing SelfDefend for GPT-3.5/4, we carefully design $P_{direct}$ and $P_{intent}$ for SelfDefend and evaluate their effectiveness in this section. As mentioned in §4, $P_{direct}$ and $P_{intent}$ represent two different defense schemes. $P_{direct}$ is more intuitive, aiming to directly pinpoint the harmful part from the original jailbreak prompt; otherwise, it outputs a token of “No”. $P_{intent}$, on the other hand, aims to recognize the true intent of the request text rather than directly intercepting harmful parts, given that indirect jailbreak attacks (e.g., DrAttack [33]) hide harmful instructions.
Based on these two guidelines, we carefully design the prompt templates used by $P_{direct}$ and $P_{intent}$. As shown in Figure 2, $P_{direct}$ uses the sentence like “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion. Inspired by chain-of-thought (CoT) [71], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then, further extracts any safety-violation part from the summarized intention, if present. It also includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning [7].
Through the extensive measurements in this section, these two designed prompts are shown to be effective within the SelfDefend framework. Nevertheless, we note that the detailed wording of these two prompt contents might have alternatives, but the current form is used as a standard template for $P_{direct}$ and $P_{intent}$ throughout the paper.
@@ 5.2 Datasets and Attack Setup
After SelfDefend is set up, we collect benchmark datasets, implement the corresponding attack methods, and determine the effectiveness metrics.
Benchmarks. Based on the five categories of existing jailbreak attacks we surveyed in §2 — human-based, optimization-based, generation-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect four benchmark datasets, JailbreakHub [62], JailbreakBench [11], MultiJail [17], and AlpacaEval [34], from which we use their user prompts for testing SelfDefend. Table 2 lists the details of our collected benchmark datasets. Specifically, we use a set of 100 harmful instructions from JailbreakBench [11], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [99] and AutoDAN [40]), generation-based jailbreaks (PAIR [12] and TAP [47]), and indirect jailbreak attacks (DrAttack [33] and Puzzler [10]). In contrast, we directly use the original prompts from the JailbreakHub, MultiJail and AlpacaEval datasets to construct inputs for the scenarios of human-based attacks, multilingual jailbreaks and normal prompts, respectively.
@@
JailbreakBench [11] 100 GCG [99], AutoDAN [40]
PAIR [12], TAP [47]
@@
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.256 0.560 0.900 0.720 0.670 0.780 0.980 0.393 0.977
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.770 0.840 1.000 0.980 0.879 0.977
GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.610 0.780 0.960 0.368 0.957
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.370 0.020 0.860 0.220 0.520 0.992
GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.290 0.020 0.710 0.220 0.203 0.972
Reduction factor 97.26% 66.07% 65.55% 59.72% 97.01% 8.97% 77.55% 48.34% 0.51%
GPT-4 (baseline) 0.047 0.080 0.190 0.330 0.310 0.740 0.900 0.076 0.973
GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.110 0.100 0.050 0.270 0.142 0.968
GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.100 0.080 0.040 0.260 0.012 0.946
Reduction factor 95.74% 100.00% 94.73% 69.69% 74.19% 94.59% 71.11% 84.21% 2.77%
GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.210 0.200 0.130 0.360 0.304 0.995
GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.180 0.170 0.130 0.280 0.019 0.970
(a) The CDF plot of $\Delta d$ for normal prompts.
(b) Average $\Delta d$ for various jailbreak prompts.
Figure 3: The extra delay $\Delta d$ caused by SelfDefend for normal and jailbreak prompts, respectively.
@@
Attack Setup. For DAN, we randomly select 1,000 samples as jailbreak queries from the forbidden question set equipped with jailbreak prompts [2], which is collected by JailbreakHub. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [13] with a batch size of 512 and 500 optimization steps. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For PAIR and TAP, we select Vicuna-13b-v1.5 [13] as the attack model. We set the maximum depth, the maximum width, and the branching factor of TAP to 10, 10, and 1, respectively. The target model of PAIR and TAP is GPT-3.5/4. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
@@
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [99] or unsafe rate [17]. Following [99, 40], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, like “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, it indicates that the jailbreak is successful. We use the common practice [99] for a list of such keywords; see Appendix B. DAN, GCG, AutoDAN, PAIR, TAP, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such cases, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal [17]. We adopt the unsafe rate in a way similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.

### v3

@@ Abstract
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance (in detection state) to concurrently protect the target LLM instance (in normal answering state) in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs can identify harmful prompts or intentions in user queries, which we empirically validate using mainstream GPT-3.5/4 models against major jailbreak attacks. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. When deployed to protect GPT-3.5/4, Claude, Llama-2-7b/13b, and Mistral, these models outperform seven state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. Further experiments show that the tuned models are robust to adaptive jailbreaks and prompt injections.
Warning: This paper contains unfiltered and potentially harmful content.
@@ 1 Introduction
The effectiveness of SelfDefend builds upon our observation that existing LLMs can identify harmful portions (prompts/intentions) in user queries, enabling the simultaneous activation of an $LLM_{defense}$ instance. To validate this hypothesis, we conduct an empirical measurement in §5 using mainstream GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [63, 102, 41, 13, 48, 34, 11, 18]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97% to 97.26% (average: 65.70%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.236, and GPT-4-based SelfDefend even reduces the ASR by 69.69% to 100% (average: 88.43%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.050. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [35] and find that the pass rate is almost unaffected for GPT-3.5 and slightly decreases by 2.77% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations. For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.06 seconds for GPT-3.5 and 0.35 seconds for GPT-4, respectively.
@@
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.050, GPT-4 itself is commonly known to be expensive [3]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [20] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data, which is then used to tune our defense models through LoRA fine-tuning [23] on the publicly available Llama-2-7b model [69]. The details will be introduced in §6.
@@
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with seven representative jailbreak defenses: ICD [75], SafeDecoding [84], Perplexity Filter [27], SmoothLLM [60], Llama Guard [26], and Llama Guard 2/3 [67, 19]. The results show that SelfDefend consistently outperforms all other defenses in 55 out of 60 tested attack scenarios (involving 10 jailbreak methods and six target LLMs: GPT-3.5, GPT-4, Llama-2-7b-chat, Mistral-7B-Instruct-v0.2, Claude-3.5-sonnet, and Llama-2-13b-chat) and reaches the defense level of GPT-4-based SelfDefend.
@@
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.39 seconds, with $\Delta d$ in all attack scenarios now below 0.1 seconds except for DAN and LLM-Fuzzer. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it efficient for potential real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [59], and empirically show that the tuned models are robust to adaptive attacks and prompt injections [45, 43]. Details are available in §7.
@@ 2.2 Jailbreak Attacks
Optimization-based Jailbreak typically updates the adversarial prompt iteratively using gradient-based or search-based methodologies [102, 64, 5, 41, 29]. The pioneering work by GCG [102] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [64] further extended this technique to GCG++ by employing a proxy model to direct the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [5] utilized a simple random search on a suffix to increase the likelihood of hitting the target probability. Unlike optimizing these obviously unreadable suffixes, AutoDAN [41] automatically constructs human-readable jailbreak prompts using a carefully designed hierarchical genetic algorithm. Furthermore, RLbreaker [14] trains a reinforcement learning agent to guide the search for adversarial prompts, making it more efficient than the stochastic mutations of JSAA and AutoDAN.
Generation-based Jailbreak employs language models [55, 17, 13, 48, 54] to generate effective jailbreak prompts that can mislead LLMs into producing restricted content. An intuitive approach is to use an auxiliary LLM to construct candidate prompts through prompt engineering. For example, PAIR [13] fed the response of the target model back to the attacking LLM to adapt the output for deceptive jailbreaks. Mehrotra et al. [48] then refined PAIR’s approach through tree-of-thought reasoning [86]. LLM-Fuzzer [88] automates jailbreak template generation for LLMs by starting with human-written templates and applying random mutations to create new inputs with the assistance of LLMs. Moreover, the adversary can train a new LLM specifically to attack the target model. For example, Paulus et al. [54] fine-tuned the Advprompter LLM to generate human-readable suffixes against the target LLM.
@@
Gradient Cuff [24] NeurIPS 2024 (2403.00867) Compare the gradient norm of refusal loss with that in benign queries
Circuit Breaking [101] NeurIPS 2024 (2406.04313) Interrupt the model output harmful content at the internal representations
@@
RPO [99], DPP [83] NeurIPS 2024 (2401.17263) Optimize universal and transferable suffixes that enforce safe outputs
@@
CAT [79] NeurIPS 2024 (2405.15589) Conduct adversarial training on continuous embedding attacks
LED [94] EMNLP 2024 (2405.18166) Identify safety-critical layers and realign them through model editing
@@ 4 The SelfDefend Framework
After analyzing the pitfalls of existing defenses, we propose a new perspective on defending LLMs against jailbreaks. Our key idea is to deploy a dedicated LLM alongside the target LLM to concurrently detect potential jailbreak queries. This idea is made possible because we found that LLMs can protect themselves by identifying harmful portions in user queries.
Insight. Normally, an LLM operates in the answering state to follow a user prompt query $P_{query}$ and return the corresponding answer response $A_{response}$. To ensure the safety of $A_{response}$, existing guardrail approaches such as Llama Guard [26] and LLM Self Defense [56] employ a model or a system prompt to assess the harmfulness of $A_{response}$ and filter it if it violates safety policies. Such an approach requires waiting for $A_{response}$ to be generated by the LLM. Our insight is that a target LLM could operate not only in the answering state but also in the detection state simultaneously, as long as we create two instances of the target model. Therefore, given the same $P_{query}$, we aim to initialize two states of the LLM at the same time, one still answering $P_{query}$ normally but the other cautiously checking $P_{query}$ (instead of answering it). This is a new perspective that has never been explored by previous works. Indeed, our measurements in §5 empirically demonstrate a significant discrepancy between the answering state and the detection state for the same LLM, with the median ASR gap reaching 2.29$\times$ for GPT-3.5 and 8.00$\times$ for GPT-4, as shown in the two rows labeled “Gap” in Table 3.
Overview. Based on this insight, we propose a generic LLM jailbreak defense framework called SelfDefend. As shown in Figure 1, SelfDefend creatively establishes a shadow stack alongside the normal stack in the LLM space to conduct checkpoint-based access control, which mimics traditional security defense concepts such as the shadow stack for defending against buffer overflow attacks [9] and the library-based checkpoint from SCLib [76]. We denote the target LLM in the normal stack as $LLM_{target}$ and the defense LLM in the shadow stack as $LLM_{defense}$. SelfDefend simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate. In SelfDefend, $LLM_{defense}$ can be instantiated from the same model as $LLM_{target}$, although in practice we suggest using a dedicated $LLM_{defense}$ that is robust and low-cost for detecting jailbreak queries.
Workflow. ① Given an incoming prompt query $P_{query}$, SelfDefend dispatches it to both $LLM_{target}$ and $LLM_{defense}$ for concurrent processing. ② $LLM_{target}$ processes $P_{query}$ as usual, whether it is a normal prompt or an adversarial prompt, but caches its token-by-token output until a checkpoint is triggered from the shadow stack. By contrast, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful part (via $P_{direct}$) or intention (via $P_{intent}$). ③ Once a token of “No” (indicating no issue) is output from the shadow stack, $LLM_{target}$ is triggered to release its token-by-token response. Otherwise, when the shadow stack detects a harmful portion (prompt/intention), SelfDefend would respond with a template that refuses to answer, i.e., “I can’t fulfill your query because your [harmful portion] violated our safety policy.” It is noted that “[harmful portion]” is replaced with the recognized portion from $LLM_{defense}$.
@@ 5 An Empirical Measurement
Besides initializing SelfDefend for GPT-3.5/4, we carefully design $P_{direct}$ and $P_{intent}$ for SelfDefend and evaluate their effectiveness in this section. As mentioned in §4, $P_{direct}$ and $P_{intent}$ represent two different defense schemes. One directly identifies the harmful part from the original jailbreak prompt, while the other indirectly recognizes the true intent of the request text. Based on these two guidelines, we carefully design the prompt templates of $P_{direct}$ and $P_{intent}$ as follows:
The design of $P_{direct}$ is based on the fact that many jailbreak attacks explicitly contain instructions to make LLMs respond with corresponding harmful content. Benefiting from the powerful semantic understanding ability of LLMs, we design $P_{direct}$ with the sentence “identify which part of the following sentences might violate your safety policies or have the jailbreak intentions” to directly pinpoint the harmful portion in the query text, as shown in Figure 2. The excerpted harmful portion indicates that the request is a jailbreak. Otherwise, $LLM_{defense}$ outputs a token of “No” when the query is safe.
The design of $P_{intent}$, on the other hand, considers that indirect jailbreak attacks (e.g., DrAttack [34]) may not explicitly contain harmful instructions. As such, $P_{intent}$ is designed to understand the true intention of the query and determine whether it contains any safety-violating content, as shown in Figure 2. Inspired by chain-of-thought (CoT) [73], $P_{intent}$ first summarizes the true intention of $P_{query}$ and then extracts any safety-violating part from the summarized intention, if present. If the query is safe, $LLM_{defense}$ also outputs “No”. Moreover, $P_{intent}$ includes two contrasting examples in the prompts to help $LLM_{defense}$ better understand the instructions, motivated by few-shot in-context learning (ICL) [8]. Both COT and ICL learning techniques have been shown to be effective in enhancing the reasoning ability of LLMs [73, 8].
Measurements conducted in this section demonstrate the effectiveness of these two prompts. While alternative wordings may exist, the current formulations for $P_{direct}$ and $P_{intent}$ serve as standardized templates throughout this paper.
@@ 5.2 Datasets and Attack Setup
Benchmarks. Based on the five categories of existing jailbreak attacks we surveyed in §2 — human-based, optimization-based, generation-based, indirect, and multilingual jailbreaks — we identify representative jailbreak attack methods in each category. We then collect four benchmark datasets, JailbreakHub [63], JailbreakBench [12], MultiJail [18], and AlpacaEval [35], from which we use their user prompts for testing SelfDefend. Table 2 lists the details of our collected benchmark datasets. Specifically, we use a set of 100 harmful instructions from JailbreakBench [12], a standardized evaluation framework, to drive optimization-based jailbreaks (GCG [102], AutoDAN [41], and RLbreaker [14]), generation-based jailbreaks (PAIR [13], TAP [48], and LLM-Fuzzer [88]), and indirect jailbreak attacks (DrAttack [34] and Puzzler [11]). In contrast, we directly use the original prompts from the JailbreakHub, MultiJail and AlpacaEval datasets to construct inputs for the scenarios of human-based attacks, multilingual jailbreaks and normal prompts, respectively.
@@
JailbreakBench [12] 100 GCG [102], AutoDAN [41], RLbreaker [14]
PAIR [13], TAP [48], LLM-Fuzzer [88]
@@
DAN GCG AutoDAN RLbreaker PAIR TAP LLM-Fuzzer DrAttack Puzzler MultiJail AlpacaEval
GPT-3.5 (baseline) 0.256 0.560 0.900 0.650 0.720 0.670 0.640 0.780 0.980 0.393 0.977
GPT-3.5-based Shadow Stack ($P_{direct}$) 0.982 0.720 0.960 0.910 0.770 0.840 0.790 1.000 0.980 0.879 0.977
GPT-3.5-based SelfDefend ($P_{direct}$) 0.242 0.450 0.870 0.600 0.600 0.610 0.500 0.780 0.960 0.368 0.957
GPT-3.5-based Shadow Stack ($P_{intent}$) 0.015 0.280 0.350 0.310 0.370 0.020 0.280 0.860 0.220 0.520 0.992
Gap between Normal and Shadow ($P_{intent}$) 17.07$\times$ 2.00$\times$ 2.57$\times$ 2.10$\times$ 1.95$\times$ 33.50$\times$ 2.29$\times$ 0.91$\times$ 4.45$\times$ 0.76$\times$ 0.98$\times$
GPT-3.5-based SelfDefend ($P_{intent}$) 0.007 0.190 0.310 0.240 0.290 0.020 0.170 0.710 0.220 0.203 0.972
Reduction factor 97.26% 66.07% 65.55% 63.08% 59.72% 97.01% 73.44% 8.97% 77.55% 48.34% 0.51%
GPT-4 (baseline) 0.047 0.080 0.190 0.290 0.330 0.310 0.190 0.740 0.900 0.076 0.973
GPT-4-based Shadow Stack ($P_{direct}$) 0.004 0.010 0.010 0.000 0.110 0.100 0.010 0.050 0.270 0.142 0.968
Gap between Normal and Shadow ($P_{direct}$) 11.75$\times$ 8.00$\times$ 19.00$\times$ $\infty$×$\times$ 3.00$\times$ 3.10$\times$ 19.00$\times$ 14.80$\times$ 3.33$\times$ 0.54$\times$ 1.01$\times$
GPT-4-based SelfDefend ($P_{direct}$) 0.002 0.000 0.010 0.000 0.100 0.080 0.000 0.040 0.260 0.012 0.946
Reduction factor 95.74% 100.00% 94.73% 100.00% 69.69% 74.19% 100.00% 94.59% 71.11% 84.21% 2.77%
GPT-4-based Shadow Stack ($P_{intent}$) 0.019 0.080 0.070 0.050 0.210 0.200 0.050 0.130 0.360 0.304 0.995
GPT-4-based SelfDefend ($P_{intent}$) 0.005 0.050 0.070 0.000 0.180 0.170 0.040 0.130 0.280 0.019 0.970
@@
Attack Setup. For DAN, we randomly select 1,000 samples as jailbreak queries from the forbidden question set equipped with jailbreak prompts [2], which is collected by JailbreakHub. For GCG, we choose its individual version and optimize the suffix on Vicuna-7b-v1.3 [15] with a batch size of 512 and 500 optimization steps. For AutoDAN, we choose its first version of the genetic algorithm, i.e., AutoDAN-GA. The genetic algorithm used in AutoDAN-GA features a crossover rate of 0.5, a mutation rate of 0.01, and 100 optimization steps. For RLbreaker, we employ GPT-3.5 as the auxiliary model to perform mutations. The total number of queries to the target LLM is limited to 10,000 for both the training and testing phases. For PAIR and TAP, we select Vicuna-13b-v1.5 [15] as the attack model. We set the maximum depth, the maximum width, and the branching factor of TAP to 10, 10, and 1, respectively. The target model of PAIR and TAP is GPT-3.5/4. For LLM-Fuzzer, we select GPT-3.5 as the helper model for mutations and set the maximum number of queries to the target LLMs at 1,000. For DrAttack and Puzzler, we use GPT-4 to construct their jailbreak prompts. For MultiJail, we selected all 315 queries in the Bengali language. For AlpacaEval, we select all 805 questions from the AlpacaEval dataset.
@@
Metrics. We measure the defense effectiveness of SelfDefend indirectly using the attack success rate (ASR) [102] or unsafe rate [18]. Following [102, 41], ASR is measured by detecting whether the LLM response contains keywords indicating a refusal to answer, such as “I can’t assist.” If such a keyword or its variant is included, it indicates that the attack is unsuccessful; otherwise, vice versa. We use the common practice [102] for a list of such keywords; see Appendix D. DAN, GCG, AutoDAN, RLbreaker, PAIR, TAP, LLM-Fuzzer, DrAttack, and Puzzler all adopt ASR as the metric for measuring jailbreak performance. The lower the ASR, the stronger the defense performance of SelfDefend against jailbreaks. Note that we also use ASR to evaluate the performance of SelfDefend under normal prompts from AlpacaEval, but a higher ASR indicates that our framework is more compatible with normal prompts (in such cases, ASR can be interpreted as answer success rate). On the other hand, the unsafe rate uses GPT-4 to determine whether the response of the target model matches the jailbreak goal [18]. We adopt the unsafe rate in a manner similar to ASR to specifically measure the jailbreak performance of MultiJail, since the response to a multilingual prompt does not contain English keywords.